find the final balance in a account after investing in principle P and letting it grow a a rate (r) for t years.
p=12,000 r=6.9% t=4 2/3 years
how do i found out the balance of $?

Answer by Theo(3464)

p = 12,000
r = .069
t = 4 + 2/3 years = 4.666666667 years
f = future value
formula:
f = p*(1+r)^t
substitute to get:
f = 12,000 * (1.069)^4.66666667
use your calculator to get:
f = 16,383.6453 which you can round up to 16,383.65.